OpenStudy (anonymous):

If f(x) = 3x + 3, g(x) = 3x + 3, and h(x) = x + 3, what is [g o f o h](–5)?

OpenStudy (anonymous):

@bohotness

OpenStudy (loser66):

open one by one. \(if~~\color{red}{ h(x) = x+3} , then f\color{red}{(h(x)}= 3(\color{red}{h(x)})+3\) \(=3(\color{red}{x+3})+3= 3x+12\) Now, you imitate to open the last one, which is g(f(h(x)) = ??? After getting the final equation, replace x =- 5 to get the answer